A car is traveling at a velocity of 37 m/s west. The car slows downs at a constant rate to a complete stop over 4 seconds. What was the acceleration of the car in m/s2?


1
Expert's answer
2021-12-13T16:55:54-0500

The acceleration is


"a=\Delta v/t=(v_f-v_i)/t=(37-0)/4=9.25\text{ m/s}^2."